2C4H18O4+13O2=8CO2+18H2O

how many atoms of oxygen are required to generate 10.3 g of carbon dioxide?
Answer #1

Molar mass of CO2,

MM = 1*MM(C) + 2*MM(O)

= 1*12.01 + 2*16.0

= 44.01 g/mol

mass of CO2 = 10.3 g

mol of CO2 = (mass)/(molar mass)

= 10.3/44.01

= 0.234 mol

According to balanced equation

mol of O2 formed = (13/8)* moles of CO2

= (13/8)*0.234

= 0.3803 mol

Mol of O atoms = 2*mol of O2

= 2*0.3803 mol

= 0.7606 mol

use:

number of atoms = number of moles * Avogadro’s number

= 0.7606 * 6.022*10^23 atoms

= 4.58*10^23 atoms

Answer: 4.58*10^23 atoms